OpenStudy (anonymous):

Please help! Suppose a triangle has two sides of length 3 and 4 and that the angle between these two sides is pi/3. What is the length of the third side of the triangle?

OpenStudy (anonymous):

\[A.\sqrt{13}\] \[B.3\] C\[\sqrt{3}\] D.\[4\sqrt{3}\]

OpenStudy (raden):

use the law of cosine

OpenStudy (raden):



OpenStudy (raden):

x^2 = 3^2 + 4^2 - 2(3)(4)cos(pi/3) solve for x (assumed the length of the third side is x)